The radius of a cylindrical gift box is ​(3x+4​) inches. The height of the gift box is twice the radius. What is the surface area of the​ cylinder? Write your answer as a polynomial in standard form.
Answer by htmentor(1343)

The surface area of a cylinder A = 2*pi*r*h (lateral) + 2*pi*r^2 (end caps), where r is the radius and h is the height.
Given: r = 3x+4; h = 2*r -> h = 2(3x+4)
A = 2*pi(2(3x+4)^2 + (3x+4)^2) = 2*pi*(3(3x+4)^2) = 6*pi*(3x+4)^2
I leave it as an exercise for you to expand and express in standard form.